OpenStudy (anonymous):

if one root of 5x^2+13+k is reciprocal of other. find k.

OpenStudy (anonymous):

a*1/a=k/5 k=5

OpenStudy (anonymous):

how it is possible

OpenStudy (anonymous):

the product of the roots of a quadratic equation is c/a where the equation is ax^2+bx+c=0 here a=5,b=13 and c=k

OpenStudy (anonymous):

how b=13 it is constant

OpenStudy (anonymous):

it must be 13x

OpenStudy (anonymous):

otherwise x^2=-(13+k)/5 if x^2=a^2 then x=a,-a not reciprocal...